Solve the equation:                       (12x-1)(6x-1)(4x-1)(3x-1) = 5

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Given eqn can be written as
(x-1/12)(x-1/6)(x-1/4)(x-1/3)=(5/12*4*6*3)..................(1)
let y=1/4[(x-1/12)+(x-1/6)+(x-1/4)+(x-1/3)]
so y=x-5/24...............(2)
Substituting 2 in 1 we get
(y+1/24)(y-1/24)(y+3/24)(y-3/24)=(5/12*4*6*3)
after simplifying and putting y2-5/242=t we get
(t+4/242)(t-4/242)=(121/1442)
t2=(16/244)+(121/1442)
t=11/144
y=\pm7/24
we get x=1/2,-1/12
Thanks and Regards
Shaik Aasif
askIItians faculty